please solve this

media%2Fd15%2Fd15d608b-6b46-4d39-bac0-eb

please solve this
Answer #1

we have given that,

Species Raccoon Bat Fox Others
probability (pi) 0.57 0.15 0.11 X

a)

we know that,

Pi=1 0.57 + 0.15 + 0.11 + X = 1 0.83+X1 X = 1-0.83 X = 0.17

b) probability of rabies is not raccoon.

p(raccoon)=0.57

p(not raccoon)=1- p(raccoon)

p(not raccoon)=1-0.57=0.43

c)

p(Bat or Fox)=p(Bat)+p(Fox)

p(Bat or Fox)=0.15+0.11

p(Bat or Fox)=0.26